### v1

# Neural Ordinary Differential Equations for Learning and Extrapolating System Dynamics Across Bifurcations

Abstract
footnotetext: Preprint
Forecasting system behaviour near and across bifurcations is crucial for identifying potential shifts in dynamical systems. While machine learning has recently been used to learn critical transitions and bifurcation structures from data, most studies remain limited as they exclusively focus on discrete-time methods and local bifurcations. To address these limitations, we use Neural Ordinary Differential Equations which provide a continuous, data-driven framework for learning system dynamics. We apply our approach to a predator-prey system that features both local and global bifurcations, presenting a challenging test case. Our results show that Neural Ordinary Differential Equations can recover underlying bifurcation structures directly from timeseries data by learning parameter-dependent vector fields. Notably, we demonstrate that Neural Ordinary Differential Equations can forecast bifurcations even beyond the parameter regions represented in the training data. We also assess the method’s performance under limited and noisy data conditions, finding that model accuracy depends more on the quality of information that can be inferred from the training data, than on the amount of data available.
1 Introduction
Bifurcation theory is a powerful tool to study how and when a system undergoes sudden shifts in its dynamical behaviour. A bifurcation is a point where a small change in the system’s parameters or external drivers cause a qualitative shift in its dynamical behaviour. Because these shifts can have significant consequences, ranging from smaller-scale ecosystem collapses [1, 2, 3, 4] to large-scale climate tipping points [5, 6, 7], understanding and predicting system behaviour near bifurcation points is crucial [8]. Identifying the types of bifurcations that a real-world system can exhibit provides insights into the system’s long-term dynamics and can help guide interventions to prevent undesired outcomes.
A major challenge lies in detecting bifurcations in systems where the underlying processes and governing equations are poorly understood or unknown. In response, data-driven approaches have gained increasing interest. In particular, machine learning is being actively explored as a tool for capturing system dynamics and bifurcations directly from data [9, 10]. Deep learning methods, for instance, outperform classical early warning indicators in both sensitivity and specificity when detecting pre-bifurcation signals in time series data[11, 12, 13]. Parameter-driven reservoir computing has successfully extrapolated the dynamics of bifurcating systems and, in some cases, even predicted behaviour beyond the training data [14, 15, 16, 17, 18]. More recently, unsupervised learning techniques, which do not require prior knowledge of the bifurcation parameter, have been shown to infer the parameter’s value directly from data [19]. Additionally, hybrid modelling, which combines mechanistic differential equations with data-driven approximators, such as neural networks, has proven to effectively capture the bifurcation behaviour of diverse systems[20, 21].
Despite these advances, most existing machine learning applications to bifurcations remain limited. Many focus exclusively on discrete machine learning methods [11, 12, 13, 19, 14, 15, 16, 18], which rely on discrete-time jumps to capture dynamics. In contrast, many real-world systems evolve smoothly with gradual shifts rather than discrete steps. This mismatch between the discrete assumptions of most machine learning methods and the continuous nature of many systems limits their effectiveness in detecting bifurcations in continuous-time settings. Moreover, much of the focus has been on local bifurcations [11, 12, 20, 21], which can be identified from local properties of the system. In contrast, global bifurcations depend on the structure of the large-scale state space. Because global bifurcations can have significant consequences for system dynamics[22], it is important to investigate methods that can infer global system structure from limited data and thereby detect this class of bifurcations.
To address these limitations, we use Neural Ordinary Differential Equations (ODEs), a class of models introduced by Chen et al.[23] that use neural networks to represent continuous-time dynamics, serving as a data-driven alternative to traditional ODEs. Instead of relying on explicitly defined equations, Neural ODEs can be used to learn a continuous vector field directly from data. This makes them well-suited for modelling complex, non-linear, or unknown systems[23, 24, 25, 26]. The continuous vector fields learned by Neural ODEs offer a promising approach for uncovering bifurcation structures from data, as they can capture smooth changes in the state space that may lead to qualitative shifts in system behaviour [24, 27].
In this work, we demonstrate the potential of Neural ODEs as a tool for extrapolating bifurcation dynamics from data. Although Neural ODEs, and even parameterized modifications[28], are established techniques, our approach introduces a novel application: explicitly incorporating a bifurcation parameter as an input to learn bifurcation structures. This specific adaptation enables the model to capture and analyse regime shifts, offering valuable insights into system behaviour. By directly learning the underlying vector field, Neural ODEs provide a global continuous-time representation of the system dynamics, making them particularly well-suited for capturing not only local bifurcations, but also global ones. We study a predator-prey system, that exhibits both local and global bifurcations, making it a challenging test case. We investigate the ability of Neural ODEs to extrapolate dynamics across bifurcations. To examine their robustness, we conduct experiments with limited trajectory data and added measurement noise, demonstrating their potential as a tool for predicting system transitions.
The rest of the paper is organized as follows. In Sec.  2, we give a brief overview on Neural ODEs and how they can be adapted to include bifurcation parameters. In Sec. 3, we introduce our predator-prey test case and summarise its bifurcation structure. In Sec. 4, we discuss the set-up of our simulation experiments. In Sec.  5, we showcase our findings, which will be further discussed in Sec. 6.
2 Neural ODEs with bifurcation parameter
Neural ODEs, introduced by Chen et al.[23], provide a framework for learning and describing continuous time dynamics using neural networks. Like classical ODEs they model system evolution through a differential equation:
$\frac{dz}{dt}=f_{\theta}(z,t),$ (1)
where $z\in\mathbb{R}^{n}$ is a vector of state variables, $f_{\theta}$ is a function represented by a neural network with $\theta$ denoting its trainable parameters (i.e., weights and biases).
Throughout this paper, we use the term model specifically to refer to the Neural ODE architecture, and model parameters to refer to its internal weights and biases. We will use the term system when referring to a dynamical system and its corresponding equations.
We use an adapted parameterized framework[28] that allows us to include bifurcation parameters. This way, we can model systems influenced by both internal state variables and external drivers. More specifically, we assume that the dynamics of the system can be described by differential equations of the form:
$\frac{dz}{dt}=f_{\theta}(z,\alpha),$ (2)
where $\alpha\in\mathbb{R}^{m}$ are the bifurcation parameters or external drivers. Note that in our formulation we omit $t$ as input, since we assume for simplicity that our system is autonomous.
The network takes the current state $z$ and the bifurcation parameter $\alpha$ as inputs and outputs the state’s time derivative. By integrating these derivatives using an ODE solver, Neural ODEs can generate system trajectories from initial conditions, combining the interpretability of traditional ODE modelling with a data-driven approach.
Representing system dynamics as a vector field allows us to simulate how the system evolves at any point in time, from any initial condition. In contrast, commonly used deep learning methods for time series forecasting, such as Recurrent Neural Networks, learn discrete-time mappings from one time point to the next, which limits them to fixed time intervals[24, 29, 30]. A continuous representation of dynamics, as used in Neural ODEs, allows us to analyse qualitative system behaviour, such as bifurcations, using approaches similar to those used in the analysis of classical ODEs.
Figure 1: Computing trajectories using a bifurcation parameter-dependent Neural ODE. Top: Zoomed in forward step. The value of the bifurcation parameter and the current state variable are used as input for the neural network. The derivative that is outputted by the neural network together with the previous state is used as input for the ODE solver, which calculates the next state. Bottom: Starting from an initial point ($z_{t_{0}}$), the Neural ODE iteratively computes the state at next time points using forward steps. The solver can take intermediate time steps between data points. The loss is only calculated at the data points.
Neural ODEs compute trajectories by iteratively integrating the learned dynamics forward in time (Fig. 1). At each forward step, the network receives the current state $z_{t^{*}}$ and bifurcation parameter $\alpha$ as inputs and outputs the corresponding time derivative of the state $\frac{dz}{dt}$. The outputted derivative together with the current state is integrated using an ODE solver, which then produces the state variables at the next time point $z_{t^{*}+\delta t}$. This iterative process allows Neural ODEs to generate trajectories that capture the underlying system dynamics. Note that we are not limited to fixed time steps, since the ODE solver can take arbitrarily small or adaptive steps to closely approximate the continuous dynamics of the system.
During training, we optimize the model parameters $\theta$ based on time series data from our system for different values of $\alpha$. During each step in the optimisation process, trajectories are generated from initial conditions, as described above. The loss is computed by comparing the predicted and true states at the observed data points. The model parameters are then updated using backpropagation, which requires computing the gradient of the loss with respect to each of the model parameters.
3 Test case: predator-prey system with local and global bifurcation
To test the bifurcation parameter-dependent Neural ODE, we applied it to a non-linear predator-prey system. In particular, the interaction between prey and predator is modelled by the following:
$\begin{split}\frac{dx}{dt}&=3x(1-x)-xy-\alpha(1-e^{-5x}),\\ \frac{dy}{dt}&=-y+3xy,\end{split}$ (3)
where $x$ and $y$ represent the prey and predator populations, respectively, and $\alpha\in\mathbb{R}^{1}$ serves as an external driver and bifurcation parameter. This system of equations represents a modification of the classic Lotka–Volterra equations with logistic growth for the prey population. The final term in Eq. 3 represents a non-linear damping effect on the prey population. This term increases with both the prey population size $x$ and the bifurcation parameter $\alpha$, indicating that higher values of $x$ and $\alpha$ lead to a larger reduction in the prey’s growth rate.
This system presents an ideal test case since the non-linear and non-polynomial dampen term allow us to evaluate the ability of the neural ODE to handle more complex dynamics. Additionally, in the parameter region we study, the system exhibits two distinct types of bifurcations: a local Hopf bifurcation and a global homoclinic bifurcation. Successfully capturing both types of bifurcations requires the model to infer both local stability changes and global transitions in state-space behaviour.
[…]
Figure 3: From trajectory data to bifurcation diagram. (a) Trajectory data is used as input to train the neural ODE. (b) The Neural ODE defines a learned vector field, which can be directly extracted from the model. (c) We determine the bifurcation structure from the vector field (Appendix C).
Fig. 2 illustrates the bifurcation structure of the system, highlighting the different dynamical regimes. Note that we limit ourself to the bifurcation parameter range $\alpha\in(0.6,0.8)$, since the zero steady state is only stable for $\alpha>0.6$. Within this range the system exhibits three distinct dynamical regimes, which allows us to analyse if the model is able to extrapolate transitions between qualitatively different behaviours. The system has four steady states: the zero steady state, two saddle nodes on the $x$-axis and a coexistence steady state. In Regime A the positive steady state is stable, meaning that for certain initial conditions predator and prey populations settle into a steady coexistence. At $\alpha=0.67$, the system undergoes a Hopf bifurcation, causing the coexistence steady state to lose stability, giving rise to stable limit cycles, which characterise Regime B. These limit cycles persist until $\alpha\approx 0.71$, where the system undergoes a global heteroclinic bifurcation; the limit cycle collides with two saddle nodes, leading to a collapse. In Regime C ($\alpha>0.71)$, the only remaining stable equilibrium is the zero steady state, meaning that both populations eventually go extinct, marking a critical transition point, that has previously been interpreted as an ecological phenomenon referred to as over-exploitation[31].
Extensions of Neural ODEs have previously been applied to more basic forms of predator-prey systems, such as the classical Lotka-Volterra system [32, 33, 34]. However, these works focused on learning dynamics from time series data with fixed parameters, using trajectories within a single dynamical regime. Consequently, they did not address how system behaviour changes with a bifurcation parameter or explore predictions across qualitatively different dynamical regimes.
4 Experiment set-up
We designed a series of simulation experiments to evaluate how well the Neural ODE captures system dynamics and generalises across bifurcations. Using simulated trajectories from the predator-prey system, we investigated whether the model can reconstruct the underlying vector field and bifurcation structure from time series data (Fig. 3).
A primary dataset was used for two purposes: first, to tune the model’s hyperparameters (see Appendix B for more details), and second, to showcase that the Neural ODE can infer the full vector field and bifurcation structure from limited trajectory data. In addition to this primary experiment, we conducted targeted experiments to explore how specific factors, such as data size and measurement noise, affect the model’s ability to learn the underlying dynamics. Below, we summarize the details of the different experiments (Fig. 4). For a more detailed overview see Appendix A.
  • Primary experiment. The data for the primary experiment consisted of trajectories sampled at five different values of the bifurcation parameter $\alpha$, with two distinct initial conditions for each value, yielding a total of ten time series. Sampling was restricted to regime A and B (as previously defined in Fig. 2), enabling us to assess whether the Neural ODE could extrapolate the global bifurcation and collapse in regime C.
  • Experiment 1: Extrapolation across different regimes. We trained the Neural ODE on trajectories from specific dynamical regimes of the bifurcation diagram to assess its ability to extrapolate to unseen regimes. In one case, the training data was drawn from both regime A and B (the primary dataset). In another, training was restricted to the smaller regime B, which exhibits only limit cycles. Lastly, we trained the model exclusively on data from regime A, where only a stable focus is present.
  • Experiment 2: Limited data. We investigated how the amount of training data influences the model performance. To explore this, we created two datasets, for which we reduced the available data compared to the primary dataset, either by decreasing the number of sampled $\alpha$-values or by computing trajectories for only a single initial condition.
  • Experiment 3: Noise levels. To study the effect of noisy observations, we introduced measurement noise into the training data and assessed the model performance compared to the primary dataset. The noisy observations were generated by applying the following transformation to our primary dataset:
    $\begin{split}\hat{x}_{i}&=x_{i}+\eta_{i},\\ \hat{y}_{i}&=y_{i}+\nu_{i},\quad\quad{\eta}_{i},{\nu}_{i}\sim\mathcal{N}(0,\sigma^{2}),\end{split}$ (4)
    where $\hat{x}_{i}$, $\hat{y}_{i}$ are our noisy data points and $\sigma$ controls the noise level. We considered three cases: no noise $(\sigma=0)$, low noise $(\sigma=0.05)$, and high noise $(\sigma=0.2)$. Using these values of $\sigma$, the noisy data we generated remained positive for both $x$ and $y$.
Figure 4: Sampling locations of training data for different experiments indicated by dashed lines. Experiment 1: Data is sampled from different regimes. Experiment 2: Amount of data is limited, either by decreasing the number of sampled $\alpha$-values or by reducing the number of initial conditions from two to one (dotted lines). Experiment 3: Different levels of measurement noise are added to the data.
The neural network architecture, ODE solver settings and hyperparameters remained fixed across all experiments. Hyperparameters were selected based on tuning performed on the primary dataset (see Appendix B for more details). For our architecture we used a Multilayer Perceptron (MLP) coupled with an adaptive ODE solver (DOPRI5[35]) for integration. The number of layers and depth of each layer were varied during the tuning process to explore different model complexities. We use the torchdiffeq library [36], which provides ODE solvers implemented in PyTorch. Backpropagation can be performed either by differentiating through all intermediate solver steps or by using the adjoint sensitivity method[23], which reduces memory usage during training. We used standard backpropagation through the internals of the solver, and did not encounter memory or stability issues.
To guide the training process, we combined a mean absolute error (MAE) with a physics-informed term to obtain our loss function:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}{\|\mathbf{z}_{i}-\tilde{\mathbf{z}}_{i}\|}_{1}+\lambda\left[\frac{1}{M}\sum_{j=1}^{M}\left|\frac{d\tilde{x}}{dt}(0,y_{j})\right|+\right.\left.\frac{1}{M}\sum_{j=1}^{M}\left|\frac{d\tilde{y}}{dt}(x_{j},0)\right|\right],$
where $\tilde{\mathbf{z}}_{i}=[\tilde{x}_{i},\tilde{y}_{i}]$ are our predicted state variables, $N$ is the number of data points and $M$ indicates the number of points on which we evaluate the physics-loss. Unlike mean squared error (MSE), which disproportionately emphasizes large errors, MAE treats all deviations linearly[37, 38]. Since MSE would assign greater weight to outliers and might obscure small but meaningful errors during training, MAE is a more suitable objective function for our application.
The physics part of the loss function enforces that the axes are invariant by penalizing non-zero derivatives of the system along specific axes. Since Neural ODEs model derivatives directly, and much of our prior knowledge about physical systems is expressed through differential equations, it is natural to incorporate physics-based information into the loss function. Such physics-informed neural ODEs have been explored in previous work [39, 40, 41]. In our experiments, adding the physics-informed term improved generalisation to unseen parameter ranges by guiding the model training in regions of the vector field not covered by the data. In Equation 4, the hyperparameter $\lambda$ is the physics-loss weight and determines how much the physics-loss contributes to the total loss. Note that the physics-informed term is not a hard constraint, but a soft constraint driving the derivatives on the axes towards zero.
5 Results
Primary Experiment
For the primary experiment, to evaluate the Neural ODE’s ability to learn vector fields from trajectory data, we trained the model on the primary dataset consisting of ten time series. As shown in Fig. 5, the model is able to reconstruct vector fields from limited trajectory data, qualitatively capturing the underlying system dynamics. Notably, the model not only replicates the behaviour observed in the training trajectories but also accurately extrapolates the onset of a global bifurcation leading to system collapse. Since global bifurcations are characterized by large-scale changes in the state space that extend beyond local trajectories, this result indicates that the model has captured structural information beyond the training data.
A closer examination of Fig. 5 reveals that the Neural ODE successfully extracted several key dynamic components. First, the learned vector field exhibits two saddle nodes along the $x$-axis, positioned similarly as those in the true system. Second, along both the $x$- and $y$-axes, the component of the flow perpendicular to the respective axis is close to zero. The derivatives not reaching exactly zero can be attributed to the soft formulation of the physics-constraint. Overall, despite minor deviations, the model accurately reconstructs the qualitative behaviour of the system across different values of $\alpha$.
Figure 5: Vector fields describing learned system dynamics. Top: True system. Bottom: Dynamics learned by the Neural ODE. Solid lines are trajectories for the initial condition $(x,y)=(1,0.01)$ to exemplify the system behaviour. The vector field for $\alpha=0.75$ is extrapolated, since this value was not included in the training range.
Figure 6: Reconstructed bifurcation diagrams for the different experiments. Each subfigure shows the true dynamics (black), best-performing learned dynamics (blue) and the learned dynamics for other random seeds (grey).
Experiment 1: Extrapolation across different regimes
We further evaluated the performance of the Neural ODE in different experiments and compared the results to those obtained in the primary experiment. Fig. 6 shows the bifurcation diagrams generated by the trained models for each experiment. Details on the simulation and visualisation process can be found in the Appendix C. The figure illustrates the model’s convergence behaviour. As in Fig. 2, limit cycles are presented by plotting the minimum and maximum value. To account for stochastic variability in initialisation and training, the model was trained using five different random seeds. The best-performing seed, based on MAE relative to the true system, is shown in blue.
When trained on time series that span both the stable focus and limit cycle regimes (regimes A and B), the model consistently predicts the occurrence of a global bifurcation across all random seeds. While the collapse is correctly captured, the system does not reach exactly zero due to the soft nature of the physics-constraint. Notably, the zero steady state is also absent from the learned dynamics in the regime on which the model was trained. This can be explained by the lack of time series in the dataset that converge to zero, making it inherently difficult for the model to learn and reproduce a state that is not represented in the training data.
When trained exclusively on time series from the limit cycle regime (regime B), the model shows varied performance across different random seeds. The best-performing seed is able to predict both the global bifurcation on the right and the transition to a stable focus on the left. However, results from other seeds are less consistent, highlighting the limited coverage of the training data and the model’s resulting uncertainty when extrapolating beyond the observed regime.
When trained solely on data from the stable focus regime (regime A), the model consistently predicts a transition to limit cycle behaviour across all random seeds, but fails entirely to capture the global bifurcation. This suggests that the training data lacks sufficient information to fully reconstruct key structures of the state space, such as saddle points along the $x$-axis, which play a crucial role in shaping the bifurcation dynamics.
Experiment 2: Limited data
To assess how the amount of training data influences the model performance, for training we used datasets with limited data compared to the primary experiment. In one case we decreased the number of sampled $\alpha$-values. This reduction introduced a varied performance for the different random seeds, notably also within the training regime. While the best-performing seed still produced a qualitatively accurate reconstruction of the bifurcation structure, other seeds showed less consistent results. Importantly, however, all random seeds correctly predicted a global collapse, indicating that the model retains some robustness in identifying critical system transitions even with limited parameter diversity.
In the other case, we restricted the number of initial conditions per $\alpha$-value. Despite this limitation, the model generally succeeded in predicting limit cycle dynamics and the transition leading to collapse. However, the predicted global bifurcation occurred at a lower $\alpha$-value than in the primary experiment.
For both types of data reduction, the best-performing random seeds were able to reproduce the qualitative behaviour of the true system. Nonetheless, a visual comparison with results from the primary experiment reveals a noticeable decline in quantitative accuracy (Appendix C), suggesting that while the overall dynamics remain consistent, the precise locations and timing of bifurcations become less reliable with less training data.
Experiment 3: Noise levels
Finally, we introduced different levels of measurement noise into the training data. At low noise levels, the predicted bifurcation parameter range and amplitude of limit cycles showed considerable variation across different random seeds. Nevertheless, the best-performing seed still produced a reasonable qualitative approximation of the system dynamics. With higher noise levels, the model’s predictions for the onset of limit cycles became less precise, exhibiting larger deviations compared to the primary, noise-free experiment. Despite this increased uncertainty, the model remained capable of distinguishing between distinct dynamical regimes, successfully identifying both the emergence of limit cycles and the eventual occurrence of the global bifurcation.
6 Discussion
In this study, we demonstrated that a Neural ODE model can successfully uncover the underlying bifurcation structure of a predator-prey system training only on limited trajectory data. Results from the experiments (Fig. 4 and 5) indicate that the model is capable of forecasting global bifurcations beyond the parameter regions represented in the training data. In Experiment 1, we restricted training data to specific regimes of the bifurcation diagram. The model’s ability to extrapolate varied depending on the chosen regime. When training trajectories did not sufficiently span the state space in order for the model to capture key dynamical features, such as saddle points, the Neural ODE had limited success in predicting transitions to other regimes. In Experiment 2, we reduced the quantity of training data compared to the primary experiment either by sampling fewer parameter values or by limiting the number of initial conditions. Although the best-performing random seeds still captured the system’s qualitative behaviour, performance became less consistent across seeds, suggesting increased model uncertainty. Finally, in Experiment 3, we introduced measurement noise into the primary dataset. Despite this added disturbance, the model retained its ability to extrapolate distinct dynamical regimes, demonstrating a certain level of robustness to noisy data.
While our results highlight the potential of Neural ODEs to learn entire vector fields directly from trajectory data, they also reveal a key challenge: the quality of extrapolation depends on the information that can be inferred from the training data. When the training data spans a limited part of the state space and essential aspects, such as saddle points, can not be inferred, the model lacks the basic components to reconstruct the full bifurcation structure. This indicates that extrapolation improves when the training data offers more extensive coverage of the system’s possible dynamic behaviours.
In real-world systems, however, it is often difficult or impossible to observe the full range of system behaviour, since features like unstable fixed points, are often not represented by the data. In such cases, integrating prior knowledge or mechanistic constraints into the learning process, e.g. through physics-informed neural networks or data augmentation using synthetic data, could enhance a model’s ability to generalise to unobserved regions[42, 43, 44, 45]. We noted a similar effect in our own work: incorporating a physics-based loss led to improved model performance, since it gave the model valuable information about the state space.
In our paper, we kept the bifurcation parameter $\alpha$ constant within each trajectory to enable more interpretable analysis of the model’s extrapolation behaviour. Allowing $\alpha$ to vary dynamically would effectively place trajectories in a higher-dimensional space, making the dynamics more complex to visualise and interpret. Starting with constant parameters therefore provided a natural and informative first step before extending the approach to scenarios involving time-dependent bifurcation parameters. While our experimental setup was deliberately simplified for clarity, the method itself is not limited to this scenario. It is broadly applicable to any continuous dynamical system for which trajectory data is available. In particular, the Neural ODE framework can incorporate dynamically-changing parameters with minimal modification, making it well-suited for real-world systems where $\alpha$ is generally known but varies over time.
In this study, we used a basic Neural ODE architecture to model the system dynamics. While this basic formulation already demonstrated strong potential, a range of extensions to the Neural ODE framework have been proposed [46, 47, 27, 32, 48, 49], including Augmented Neural ODEs [50], which may offer improved accuracy in capturing complex dynamical behaviours or greater robustness in the presence of noise. A promising direction for future work would be to systematically compare the performance of standard Neural ODEs, their extensions, and discrete-time machine learning methods in tasks such as forecasting global bifurcations and reconstructing bifurcation diagrams from data. While such comparisons are not straight-forward, especially since discrete-time models typically lack access to the underlying vector field and operate at fixed time steps, they could provide valuable insight into the trade-offs between continuous and discrete approaches in terms of accuracy, interpretability, and computational cost.
While the above comparisons focus on different machine learning models for forecasting and reconstruction, another valuable perspective is to examine how the proposed approach compares to, and could potentially be integrated with, alternative frameworks for predicting critical transitions. A common approach to predicting critical transitions, specifically local bifurcations, is through early warning signals based on slowed recovery after small perturbations [51, 52, 53]. Unfortunately, these signal are not observed before all tipping points, including global bifurcations [54, 55, 56]. Furthermore, because these signals do not model or capture the system’s dynamics, they cannot be used to extrapolate behaviour beyond the bifurcation point [57]. This limitation points to a potential integration with Neural ODEs. Since early warning signals typically rely on process noise to reveal subtle changes in stability, introducing similar noise-driven variability into the training data could help the Neural ODE detect potential stability changes in trajectory data more reliably. In future work, these signals could potentially be treated as an additional input to improve the model’s ability to anticipate and reconstruct complex bifurcations.
In conclusion, this work demonstrates that Neural ODEs can be used not only to model continuous-time dynamics from trajectory data, but also to recover underlying bifurcation structures, even in the presence of global bifurcations. By explicitly including the bifurcation parameter and learning the full vector field, our approach offers a flexible and interpretable framework for studying critical transitions in dynamical systems. These results motivate future extensions to more realistic systems, including those with multiple or time-varying drivers and unknown environmental interactions.
Acknowledgements
The authors acknowledge Wageningen University and Research for their investment program Data Science and Artificial Intelligence.
Data Availability
All code and data that support the findings of this study are available on Github:
https://github.com/evantegelen/neural-bifurcations.
Appendix A Data overview
In this section, we provide additional details about the data used to train the models for the various experiments. Table 1 summarizes the initial conditions and bifurcation parameter values that were used to generate the time series data. For each configuration, we used numerical integration to simulate the solution trajectories and obtained time series data by sampling at discrete time steps $t={0,1,2,\dots,100}$. Note that in each of the experiments one of the datasets overlaps with the primary experiment. This was done for easy comparison.
Fig. 7 provides an example of the trajectory data for a specific value of $\alpha$. To highlight the impact of the added measurement noise, the figure also displays the corresponding time series under different noise levels, illustrating how the original dynamics become increasingly distorted as the noise level increases.
Table 1: Overview of experimental setup with $\alpha$-values, initial conditions, and noise levels for the different datasets. Differences with respect to the primary experiment are indicated in cursive.
Bifurcation parameter values Initial conditions Noise levels
[…]
Primary experiment
0.62, 0.64, 0.66, 0.68, 0.7 (1,0.01),(1,0.1) -
Experiment 1
A and B 0.62, 0.64, 0.66, 0.68, 0.7 (1,0.01),(1,0.1) -
Only B 0.68, 0.685, 0.69, 0.695, 0.7 (1,0.01),(1,0.1) -
Only A 0.61, 0.62, 0.63, 0.64, 0.65 (1,0.01),(1,0.1) -
Experiment 2
10 time series, 5 $\alpha$-values 0.62, 0.64, 0.66, 0.68, 0.7 (1,0.01),(1,0.1) -
6 time series, 3 $\alpha$-values 0.62, 0.66, 0.7 (1,0.01),(1,0.1) -
5 time series, 5 $\alpha$-values 0.62, 0.64, 0.66, 0.68, 0.7 (1,0.1) -
Experiment 3
No noise 0.62, 0.64, 0.66, 0.68, 0.7 (1,0.01),(1,0.1) -
Low noise 0.62, 0.64, 0.66, 0.68, 0.7 (1,0.01),(1,0.1) 0.05
High noise 0.62, 0.64, 0.66, 0.68, 0.7 (1,0.01),(1,0.1) 0.2
[b]
Figure 7: Time series data for $\alpha=0.62$ for with different noise levels.
Appendix B Hyperparameters and tuning
We tuned our model hyperparameters once for the primary experiment. To perform hyperparameter tuning, we split our primary dataset into a training and validation set, ensuring that the validation set contained a value of $\alpha$ outside the range seen during training to specifically test the model’s ability to extrapolate. More specifically, from the primary dataset we selected the two trajectories with $\alpha=0.62$ for validation, while the remaining eight trajectories were used for training.
Hyperparameter selection was conducted via a full grid search over the ranges specified in Table 2. Our aim was not to optimize model performance, but simply to demonstrate that Neural ODEs have the potential to learn bifurcation structures, so we explored only a limited range of tuning parameters. Each configuration was evaluated across three different random seeds to account for stochastic variability in model initialization and training. For each trial, we computed the following MAE on the validation set:
[…]
where $\mathbf{z}_{i}=[x_{i},y_{i}]$ are our validation data points, $\tilde{\mathbf{z}}_{i}$ our predicted state variables, and $N_{val}$ is the number of validation data points. Note that the error on the validation set is calculated differently from the loss function in Eq. 4, since the physics-loss is not taken into account. The hyperparameter combination with the lowest average MAE across the different seeds was selected and used to train the models for all experiments.
Table 2: Hyperparameters considered during tuning, along with their tested values. Best performing hyperparameter combination is indicated in bold (lowest average MAE over three different random seeds).
Hyperparameter Values
Learning rate 0.0001, 0.001
$\lambda$ (physics loss weight) 0.1, 0.01, 0.001
Number of layers 1, 2, 3
Layer depth 32, 64
In addition to the tuned hyperparameters, several other parameters were held fixed throughout all experiments. These are listed in Table 3.
Table 3: Fixed hyperparameters used throughout all experiments. These were not varied during tuning.
Hyperparameter Value
Relative ODE solver tolerance ($r_{\text{tol}}$) $1\times 10^{-5}$
Batch size 5
Batch length 20
Number of epochs 5000 (during tuning), 10000 (final experiments)
During training, we employed batching to improve convergence and efficiency. All time series were segmented into non-overlapping blocks of a fixed number of time steps, defined by the batch length. A training batch was then formed by randomly selecting a number of these blocks (equal to the batch size). For each batch, the loss was computed and the network parameters were updated accordingly. At every epoch, new batch combinations were generated by randomly re-sampling blocks from the full training set, ensuring diverse training input across epochs.
Appendix C Generating bifurcation diagrams from vector fields
To investigate the dynamical behaviour of our trained models, we used a simulation-based approach to generate the bifurcation diagrams (shown in Fig. 6). These simulations were done after training the models using the learned vector fields. For 500 different values of the bifurcation parameter $\alpha$, we ran each model until $t=2000$ from two distinct initial conditions: $(x,y)=(0,0)\nolinebreak$ and $(x,y)=(1,0.01)\nolinebreak$. From the final 500 time points of each simulation, we extracted the minimum and maximum value for the $x$-component, which serve as indicators of the system’s long-term behaviour. If the model converges to a fixed point, these values coincide; in the case of a limit cycle, they reflect the oscillation amplitude in the $x$-direction. To ensure a fair comparison, we applied the same procedure to the true system using identical initial conditions and parameter values. We restrict our bifurcation diagrams for the $x$-component, as it captures the most informative aspects of the system dynamics.
[…]
Primary experiment 0.036 $\pm$ 0.004
Experiment 1
A and B 0.036 $\pm$ 0.004
Only B 0.074 $\pm$ 0.009
Only A 0.092 $\pm$ 0.007
Experiment 2
10 time series, 5 $\alpha$-values 0.036 $\pm$ 0.004
6 time series, 3 $\alpha$-values 0.072 $\pm$ 0.014
5 time series, 5 $\alpha$-values 0.087 $\pm$ 0.025
20 time series, 10 $\alpha$-values 0.033 $\pm$ 0.006
Experiment 3
No noise 0.036 $\pm$ 0.004
Low noise 0.054 $\pm$ 0.011
High noise 0.059 $\pm$ 0.005
Table  4 shows quantitatively how the model performed for the different experiments. Similar conclusions can be drawn as from the visual representation in Fig. 4. For Experiment 1, when comparing to the primary experiment, only training on regime B decreased the overall accuracy and increased the variation between the different seeds. The accuracy decreased even further when only training on regime A.
For Experiment 2, it can be seen that decreasing the amount of data does impact the accuracy quite a bit. There is also a clear increase in the standard deviation, indicating that the differences in performance between the seeds increases. This coincides with what we saw in Sec. 5; some random seeds were still able to reconstruct the bifurcation structure qualitatively well, but others failed. The table also shows an additional dataset, where the amount of data was increased, of which the results are discussed in Sec. D. Interestingly, doubling the amount of time series data does not lead to a big decrease in the MAE.
For Experiment 3, the table shows that adding measurement noise to the data indeed impacts the performance of the model. With low noise levels the accuracy goes down compared to the primary experiment, but also the performance varies more between seeds. For high levels of noise there is less varied performance between the seeds, but the overall accuracy is lower.
Appendix D Additional experiment: increase data
We present an additional experiment in which the amount of data was increased to investigate the impact of data availability on model performance. Specifically, we doubled the number of sampled $\alpha$-values in the training set compared to the primary dataset. However, this increase did not lead to a substantial improvement in the model’s accuracy (Fig. 8). The additional samples failed to provide new information for the Neural ODE to further refine its learned dynamics. This suggests that the model could infer no new dynamics that were not yet represented in the primary dataset. These results highlight that, in this setting, simply increasing the amount of data does not necessarily translate to better generalization or accuracy.
Figure 8: Reconstructed bifurcation diagram for dataset with increased amount of data compared to the primary dataset. Left: primary dataset. Right: bifurcation diagram learned from dataset that included 20 time series for 10 different $\alpha$-values.